OpenStudy (anonymous):

Simplify. 4y-20/6-8y-12/24 Explain.

OpenStudy (anonymous):

=4y-8y-(80/24)-(12/24) =-4y-(92/24) =-4y-(23/6)

OpenStudy (anonymous):

=4y + -8y = -4y =-20/6 + -12/24 = -4y - (23/6) -23/6 can be written as -3 5/6 as well in mixed form. Explanation: You add like terms, which is 4y and -8y together, separately 20/6 and -12/24.